OpenStudy (anonymous):

What is the sum of a 7-term geometric series if the first term is 6, the last term is 24,576, and the common ratio is 4? −32,766 −19,662 19,662 32,766

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

i got -32, 766

jimthompson5910 (jim_thompson5910):

\[\Large S_n = a*\frac{1-r^n}{1-r}\] Plug in a = 6 (first term) r = 4 (common ratio) n = 7 (number of terms)

OpenStudy (anonymous):

6* 1-4^7/1-4?

jimthompson5910 (jim_thompson5910):

I think you meant this right? \[\Large 6*\frac{1-4^7}{1-4}\]

OpenStudy (anonymous):

yeah!

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

compute that with a calculator

jimthompson5910 (jim_thompson5910):

make sure you use parenthesis

OpenStudy (anonymous):

i got 32766

jimthompson5910 (jim_thompson5910):

same here

OpenStudy (anonymous):

thanks again

jimthompson5910 (jim_thompson5910):

np

OpenStudy (anonymous):

will you check this one? @jim_thompson5910

OpenStudy (anonymous):

What is the sum of the geometric sequence 1, −6, 36, … if there are 7 terms? −39,991 −6,665 6,665 39,991 x this one?

jimthompson5910 (jim_thompson5910):

Yep \[\Large S_n = a*\frac{1-r^n}{1-r}\] \[\Large S_7 = 1*\frac{1-(-6)^7}{1-(-6)}\] \[\Large S_7 = 39,991\]